geerky42 (geerky42):

\(\large b^x = b^y \Rightarrow x = y \) iff \(\large b > 0, b \neq 1\) Why b>0? Why not b ≠ 0?

OpenStudy (freckles):

(-1)^3=(-1)^5 => 3=5. True or false?

geerky42 (geerky42):

Ok, how about b≠-1, 0, 1? I just don't understand why b couldn't be smaller than zero...

OpenStudy (freckles):

-1 is smaller than 0

OpenStudy (freckles):

The statement did not work for b<0

geerky42 (geerky42):

except -1... Why couldn't it be smaller than zero?

geerky42 (geerky42):

except -1...

geerky42 (geerky42):

Why not \(b \in \mathbb{R}, b \neq -1, 0, 1\)?

OpenStudy (freckles):

Is the the thingy suppose to go both ways?

OpenStudy (freckles):

Or do you mean it just in that one way?

geerky42 (geerky42):

Well, both way, I guess.

geerky42 (geerky42):

I found this statement in a iPhone app called Math Formulas, I think this is wrong, but I'm not sure...

OpenStudy (freckles):

Well it is probably leading up to logarithms... Of course 1^n=1^m but this does not imply n=m.

OpenStudy (freckles):

Oh you understand why b cannot be -1,0, or 1.

OpenStudy (freckles):

Have you talked about logarithms?

OpenStudy (freckles):

\[\log_b(x)=\frac{\ln(x)}{\ln(b)} , x>0, b>0, b \neq 1\]

OpenStudy (freckles):

\[b^x=b^y\] \[\log_b(b^x)=\log_b(b^y)\] \[x \log_b(b)=y \log_b(b)\] \[x(1)=y(1)\] \[x=y\]

OpenStudy (freckles):

That is assuming b>0 and b does not equal 1.

geerky42 (geerky42):

Well, this makes sense. Thanks.